# Linear fractional relative risk aversion

### v1

@@ Introduction
Corollary 1 (Special cases)
Proof.
@@
[…]
$\widetilde{z}v^{\prime\prime}(\widetilde{z})+(\beta-\widetilde{z})v^{\prime}(\widetilde{z})=0,$
[…]
@@
Letting $\sigma=1$ in (8), we have
@@
[…]
$\widetilde{z}v^{\prime\prime}(\widetilde{z})+(1-\widetilde{z})v^{\prime}(\widetilde{z})=0,$
which is the confluent hypergeometric differential equation $\widetilde{z}v^{\prime\prime}(\widetilde{z})+(c-\widetilde{z})v^{\prime}(\widetilde{z})-av(\widetilde{z})=0$ with coefficients $a$ and $c$ satisfying $a=0$ and $c=1$. This can be rewritten as $\frac{\widetilde{z}v^{\prime\prime}(\widetilde{z})+v^{\prime}(\widetilde{z})}{\widetilde{z}v^{\prime}(\widetilde{z})}=1$, so that $\frac{{\rm d}\ln(\widetilde{z}v^{\prime}(\widetilde{z}))}{{\rm d}\widetilde{z}}=1$. The differential equation yields $\ln\bigl{(}\widetilde{z}v^{\prime}(\widetilde{z})\bigr{)}=\widetilde{z}+c_{1}$, and thus $v^{\prime}(\widetilde{z})={\rm e}^{c_{1}}\frac{{\rm e}^{\widetilde{z}}}{\widetilde{z}}$, where $c_{1}$ is a constant of integration. Integrating the foregoing equation yields
$\int_{c_{2}}^{\widetilde{z}}v^{\prime}(t){\rm d}t={\rm e}^{c_{1}}\int_{c_{2}}^{\widetilde{z}}\frac{{\rm e}^{t}}{t}{\rm d}t,$
@@
$\int_{c_{2}}^{\widetilde{z}}v^{\prime}(t){\rm d}t=\int_{-c_{2}}^{-\widetilde{z}}w^{\prime}(s){\rm d}s=-{\rm e}^{c_{1}}\int_{-c_{2}}^{-\widetilde{z}}\frac{{\rm e}^{-s}}{-s}{\rm d}s={\rm e}^{c_{1}}\int_{-\widetilde{z}}^{-c_{2}}\frac{{\rm e}^{-s}}{-s}{\rm d}s={\rm e}^{c_{1}}\left(\int_{-\widetilde{z}}^{\infty}\frac{{\rm e}^{-s}}{-s}{\rm d}s-\int_{-c_{2}}^{\infty}\frac{{\rm e}^{-s}}{-s}{\rm d}s\right).$
Let ${\rm Ei}(\widetilde{z})=-\int_{-\widetilde{z}}^{\infty}\frac{{\rm e}^{-s}}{s}{\rm d}s$ denote the exponential integral. We then have $w(-\widetilde{z})-w(-c_{2})=v(\widetilde{z})-v(c_{2})={\rm e}^{c_{1}}[{\rm Ei}(\widetilde{z})-{\rm Ei}(c_{2})]$, which can be rewritten as $v(\widetilde{z})={\rm e}^{c_{1}}{\rm Ei}(\widetilde{z})+v(c_{2})-{\rm e}^{c_{1}}{\rm Ei}(c_{2})$. Setting $C=v(c_{2})-{\rm e}^{c_{1}}{\rm Ei}(c_{2})$ and $K={\rm e}^{c_{1}}$, and recalling that $v(\widetilde{z})=u(q)$ and $\widetilde{z}=-\alpha q$, we obtain
@@
Table 2 summarizes the special cases discussed in Corollary 1.
Table 2: Special cases of utility functions.
HARA Hypergeometric CREMR
$\beta=0$ $\beta\in(0,1)$ $\beta=1$
$\sigma=-1$ $u(q)=-\frac{K}{2}(\alpha q^{2}-2q)$
Notes: We suppress the additive constant $C$ in each expression.
We have characterized the utility function that gives rise to the LFRRA, and we have shown how to obtain special cases. Another question is how our CREMR utility relates to that proposed by Mrázová et al. (2021). Recall that the last part of Proposition 2, where $\beta$ goes to $1$ from below, does not apply to the case with $\alpha\sigma<0$, whereas Mrázová et al. (2021) have no restriction on the sign of their $\gamma$ parameter. Yet, we can nest this case in our approach by relaxing our assumptions that $K>0$, and that $1+\alpha\sigma q>0$. Indeed, assuming that $K\in\mathbb{C}$, and that $1+\alpha\sigma q<0$, where $\mathbb{C}$ is the set of complex numbers, we can retain the positive and real-valued marginal utility in (9). By making a change of variables we can subsume the CREMR utility function in Mrázová et al. (2021, their equation (12)) as follows:
Corollary 2 (Another special case)
which requires the quantity restriction $q>\gamma$ for $\gamma>0$. [^25] Although $K$ can be complex and $\alpha\sigma<0$ when $\gamma>0$, this poses no problem because the complex term $(-\frac{1}{\gamma})^{\frac{1}{\sigma}}$ in $K$ is offset by the complex term $(1+\alpha\sigma q)^{1-\frac{1}{\sigma}}$ in (13), which equals $\bigl{(}1-\frac{q}{\gamma}\bigr{)}^{1-\frac{1}{\sigma}}=\bigl{[}\bigl{(}-\frac{1}{\gamma}\bigr{)}(q-\gamma)\bigr{]}^{1-\frac{1}{\sigma}},$ to yield a real-valued utility.
Proof.
@@
We thus obtain (20). $\square$
Note that Mrázová et al. (2021) impose $\sigma>1$ in (20). The same restriction is required for monopolistic competition that we analyze in the next section, where profit maximizing prices depend on $\sigma$. However, it is sufficient to assume that $\sigma>0$ in different contexts (e.g., when equation (20) is embedded in a perfectly competitive model with intertemporally optimizing consumers).
Monopolistic competition
There is a mass $L$ of consumers in the economy. Let ${\cal U}\equiv\int_{\Omega}u(q(\omega)){\rm d}\omega$ denote an additively separable upper-tier utility function for the representative consumer, where $u$ is a utility function satisfying both $u^{\prime}>0$ and $u^{\prime\prime}<0$, $q(\omega)\geq 0$ is consumption of variety $\omega$, and $\Omega$ denotes the set of varieties available in the economy. The representative consumer maximizes $\cal U$ subject to the budget constraint $\int_{\Omega}p(\omega)q(\omega){\rm d}\omega=y$, where $p(\omega)$ is the price of variety $\omega$ and $y$ is the income of that consumer. This yields the first-order condition $u^{\prime}(q(\omega))=\nu p(\omega)$, where $\nu>0$ is the Lagrange multiplier.
A firm with marginal cost $m$ maximizes operating profits (pm)Lqf(p-m)Lq-f, where $Lq$ is the aggregate demand for the firm’s variety. Using the first-order condition for utility maximization, the firm’s profit maximization problem is given by $\max_{q}\bigl{[}\frac{u^{\prime}(q)}{\nu}-m\bigr{]}Lq-f$. Since there is a continuum of firms, no firm can affect $\nu$. Thus, the first-order condition for profit maximization is given by $\frac{u^{\prime\prime}(q)}{\nu}Lq+\bigl{[}\frac{u^{\prime}(q)}{\nu}-m\bigr{]}L=0$, and the profit maximizing price $p(m)$ satisfies:
where $\mu(m)\equiv\frac{m}{p(m)}$ denotes the inverse of the markup. For the profit-maximizing price to be valid, we focus on the case with $p(m)=\frac{m}{\mu(m)}\in[m,\infty)$, which implies $\mu(m)\in(0,1]$.
The second-order condition (SOC) for profit maximization is given by $\frac{u^{\prime\prime\prime}(q(m))}{\nu}Lq(m)+\frac{u^{\prime\prime}(q(m))}{\nu}L+\frac{u^{\prime\prime}(q(m))}{\nu}L=\frac{u^{\prime\prime\prime}(q(m))}{\nu}Lq(m)+2\frac{u^{\prime\prime}(q(m))}{\nu}L=-\frac{Lu^{\prime\prime}(q(m))}{\nu}\bigl{[}-\frac{q(m)u^{\prime\prime\prime}(q(m))}{u^{\prime\prime}(q(m))}-2\bigr{]}<0$. Since $\frac{Lu^{\prime\prime}(q(m))}{\nu}>0$, the SOC reduces to
$2-\left[-\frac{q(m)u^{\prime\prime\prime}(q(m))}{u^{\prime\prime}(q(m))}\right]>0,$ (22)
which must hold for any $q(m)$ satisfying the first-order condition for profit maximization. In what follows, we suppress $m$ to simplify the notation when there is no confusion.
3.1 Price equilibrium
$q=\frac{1}{\alpha}\frac{1-\beta-\mu}{1-\sigma(1-\mu)}.$ (23)
Combining these two cases, we obtain the following proposition.
Proposition 3 (Ranges of markups and quantities)
Proof.
We have already analyzed the case where either $\alpha=0$ or $\beta=\frac{1}{\sigma}$ holds. Thus, in this proof, we focus on the case where $\alpha\neq 0$ and $\beta\neq\frac{1}{\sigma}$.
@@ 3.1 Price equilibrium
We first derive the ranges of markups and quantities that satisfy the four inequality conditions, $q\geq 0$, $u^{\prime}>0$, $u^{\prime\prime}<0$, and $0<\mu\leq 1$. Table 3 summarizes these ranges.
Table 3: The ranges of markups and quantities satisfying $q\geq 0$, $u^{\prime}>0$, $u^{\prime\prime}<0$, and $0<\mu\leq 1$.
@@
$\Phi(q)\equiv\left\{\begin{array}[]{ll}\alpha^{2}(\sigma-1)\bigl{[}q+\frac{1-\beta}{\alpha(\sigma-1)}\bigr{]}^{2}-\frac{(1-\beta)(1-\beta\sigma)}{\sigma-1}&\mbox{for}\ \sigma\neq 1\\ (1-\beta)(2\alpha q+\beta)&\mbox{for}\ \sigma=1\end{array}\right.$. (25)
@@
$q^{\pm}=-\frac{1-\beta\pm\Delta}{\alpha(\sigma-1)},\quad\mbox{where}\quad\Delta\equiv\sqrt{(1-\beta)(1-\beta\sigma)}.$ (26)
@@
First, assume that $\sigma<1$. Since $1-\beta\sigma>0$ and $\beta\neq 1$ hold, $\Phi(q)=0$ has two real roots (26). Consider the subcase with $\beta\in(0,1)$. In that case, one of them is positive since $1-\beta+\Delta>0$, whereas the other is negative since $1-\beta-\Delta<0$. Furthermore, $\Phi$ is concave, so that it is positive for $0\leq q<-\frac{1-\beta+\Delta}{\alpha(\sigma-1)}$. We also know from Table 3 that $q<-\frac{1-\beta}{\alpha(\sigma-1)}$ must hold. Since $\frac{1-\beta}{\alpha(\sigma-1)}<-\frac{1-\beta+\Delta}{\alpha(\sigma-1)}$ holds for $\sigma<1$, the SOC holds for all $0\leq q<-\frac{1-\beta}{\alpha(\sigma-1)}$ when $\beta\in(0,1)$. Consider the subcase with $\beta=0$. Then, the two real roots in (26) become $\frac{2}{\alpha(\sigma-1)}>0$ and 0. Since $\Phi$ is concave and $\frac{2}{\alpha(\sigma-1)}$ is greater than $\frac{1-\beta}{\alpha(\sigma-1)}$ in Table 3 evaluated at $\beta=0$, the SOC holds for all $0<q<-\frac{1}{\alpha(\sigma-1)}$. As $q$ goes to zero, we know that $\frac{\Phi(q)}{(\alpha q+\beta)(\alpha\sigma q+1)}$ evaluated at $\beta=0$ goes to 2. Hence, the SOC holds for all $0\leq q<-\frac{1-\beta}{\alpha(\sigma-1)}$ when $\beta\in[0,1)$.
Second, assume that $\sigma=1$. Then, the SOC holds for all $q\geq 0$ since both $1-\beta$ and $2\alpha q+\beta$ in (25) are positive when $\beta\in(0,1)$ and since $\lim_{\sigma\to 1}\frac{\Phi(q)}{(\alpha q+\beta)(\alpha\sigma q+1)}=\frac{2}{\alpha q+1}>0$ when $\beta=0$. Hence, when $\sigma=1$ and $\beta\in[0,1)$, the SOC holds for all $0\leq q<\infty$ given in Table 3 (recall that when $\sigma=1$, $0\leq q<-\frac{1-\beta}{\alpha(\sigma-1)}$ must be replaced with $0\leq q<\infty$).
Last, assume that $\sigma>1$. Since $1-\beta\sigma>0$ and $\beta\neq 1$ hold, $\Phi(q)=0$ has two real roots (26). Consider the subcase with $\beta\in(0,1)$. In that case, both are negative since $1-\beta+\Delta>0$ and $1-\beta-\Delta>0$. Furthermore, $\Phi$ is convex, so that it is positive for all $q\geq 0$. We also know from Table 3 that there is no additional restriction on $q$, so that the SOC holds for all $0\leq q<\infty$ when $\beta\in(0,1)$. Consider the subcase with $\beta=0$. Then, the two real roots in (26) become $\frac{2}{\alpha(\sigma-1)}<0$ and 0. Since $\Phi$ is convex, the SOC holds for all $q>0$. As $q$ goes to zero, we know that $\frac{\Phi(q)}{(\alpha q+\beta)(\alpha\sigma q+1)}$ evaluated at $\beta=0$ goes to 2. Hence, the SOC holds for all $0\leq q<\infty$ when $\beta\in[0,1)$.
Since the range of $q$ in Table 3 is unaffected by the SOC, that of $\frac{1}{\mu}$ in Table 3 is also unaffected because $\mu=1-\frac{\alpha q+\beta}{\alpha\sigma q+1}$ is monotone with respect to $q$. Hence, the ranges for markups and quantities in Table 3 remain valid in Table 4.
@@
If $\beta\in(0,1)$, then $\Phi(q)=0$ has no real roots since $1-\beta\sigma<0$. Since $\sigma>1$ and $\beta\not\in\{0,1\}$, $\Phi$ is convex and $\Phi(0)=\beta(1-\beta)>0$, which implies that $\Phi$ is positive for all $q\geq 0$. Hence, the SOC holds for all $q\geq 0$. Since the range of $q$ in Table 3 is unaffected by the SOC, that of $\frac{1}{\mu}$ in Table 3 is also unaffected because $\mu=1-\frac{\alpha q+\beta}{\alpha\sigma q+1}$ is monotone with respect to $q$.
If $\beta=1$, then $\Phi(q)=0$ yields $q=0$, which implies $\Phi(0)=0$. Since $\sigma>1$, $\Phi$ is convex, which implies that $\Phi$ is positive for all $q>0$. Hence, the SOC holds for all $q>0$. Since the range of $q$ in Table 3 is unaffected by the SOC, that of $\frac{1}{\mu}$ in Table 3 is also unaffected because $\mu=1-\frac{\alpha q+\beta}{\alpha\sigma q+1}$ is monotone with respect to $q$.
@@
First, assume that $\sigma<1$. Consider the subcase with $\beta\in(0,1)$. In that case, $\Phi(q)=0$ has two real roots (26). One of them is negative since $1-\beta+\Delta>0$, whereas the other root is positive since $1-\beta-\Delta<0$. Furthermore, $\Phi$ is concave, so that it is positive for $0\leq q<-\frac{1-\beta-\Delta}{\alpha(\sigma-1)}$. We also know from Table 3 that $0\leq q<-\frac{\beta}{\alpha}$ must hold. Since $\frac{1-\beta-\Delta}{\alpha(\sigma-1)}<-\frac{\beta}{\alpha}$ holds, the second-order condition holds for all $0\leq q<-\frac{1-\beta-\Delta}{\alpha(\sigma-1)}$ but it does not hold for $\frac{1-\beta-\Delta}{\alpha(\sigma-1)}\leq q<-\frac{\beta}{\alpha}$. Since the bounds for $q$ are modified, we need to modify those for $\mu$. Noting that $\frac{\partial\mu}{\partial q}=\frac{\partial}{\partial q}\bigl{(}1-\frac{\alpha q+\beta}{\alpha\sigma q+1}\bigr{)}=-\frac{\alpha(1-\beta\sigma)}{(\alpha\sigma q+1)^{2}}>0$, we evaluate $\mu=1-\frac{\alpha q+\beta}{\alpha\sigma q+1}$ at $q=0$ and $q=-\frac{1-\beta-\Delta}{\alpha(\sigma-1)}<-\frac{\beta}{\alpha}$ to obtain $1-\beta\leq\mu<\frac{(1-\beta)\sigma+\Delta}{1+(1-\beta)\sigma}<1$. [^27] The last inequality can be obtained as follows. First, for $\frac{(1-\beta)\sigma+\Delta}{1+(1-\beta)\sigma}<1$ to hold, we require $\frac{-1+\Delta}{1+(1-\beta)\sigma}<0$. Second, the latter inequality requires that the numerator and the denominator must have opposite signs, which is indeed the case. Hence, we obtain the bounds for markups $1<\frac{1+(1-\beta)\sigma}{(1-\beta)\sigma+\Delta}<\frac{1}{\mu}\leq\frac{1}{1-\beta}$ for $\sigma<1$. Consider the subcase with $\beta=1$. Then, $\Phi$ attains the maximum $\Phi(q)=0$ at $q=0$. Since $\sigma<1$, $\Phi$ is concave, so that the second-order condition does not hold for $q\geq 0$, thus excluding the second line in Case 2a in Table 3.

### v2

@@ Appendix A.4 Proof of Proposition 2
Appendix A.5 Proof of Corollary 1
@@
In this case, equation (1) reduces to $\frac{qu^{\prime\prime}(q)}{u^{\prime}(q)}=\alpha q+\beta$, which can be rewritten as $qu^{\prime\prime}(q)+(\alpha q+\beta)u^{\prime}(q)=0$. Let $z=-\alpha q$ and $v(z)=u(q)$, so that $u^{\prime}(q)=-\alpha v^{\prime}(z)$ and $u^{\prime\prime}(q)=(-\alpha)^{2}v^{\prime\prime}(z)$. Thus, we have $\frac{z}{\alpha}(-\alpha)^{2}v^{\prime\prime}(z)+(-z+\beta)(-\alpha)v^{\prime}(z)=0$. This implies
$zv^{\prime\prime}(z)+(\beta-z)v^{\prime}(z)=0,$
which is the confluent hypergeometric differential equation $zv^{\prime\prime}(z)+(c-z)v^{\prime}(z)-av(z)=0$ with coefficients $a$ and $c$ satisfying $a=0$ and $c=\beta$. Letting $c_{1}$ and $c_{2}$ denote arbitrary coefficients, the solution to this differential equation is given by $v(z)=c_{1}\times{}_{1}F_{1}(a;c;z)+c_{2}\times z^{1-c}{}_{1}F_{1}(1+a-c;2-c;z)$, where $_{1}F_{1}(a;c;z)=\sum_{n=0}^{\infty}\frac{(a)_{n}}{(c)_{n}}\frac{z^{n}}{n!}$, is the confluent hypergeometric function (see Seaborn, 1991). Noting that $_{1}F_{1}(0;\beta;z)=1$, we have
@@
Letting $\sigma=1$ in (A.3), we have
@@
In this case, equation (1) reduces to $\frac{qu^{\prime\prime}(q)}{u^{\prime}(q)}=\alpha q+1$, which can be rewritten as $qu^{\prime\prime}(q)+(\alpha q+1)u^{\prime}(q)=0$. Let $z=-\alpha q$ and $v(z)=u(q)$, so that $u^{\prime}(q)=-\alpha v^{\prime}(z)$ and $u^{\prime\prime}(q)=(-\alpha)^{2}v^{\prime\prime}(z)$. Thus, we have $\frac{z}{\alpha}(-\alpha)^{2}v^{\prime\prime}(z)+(-z+1)(-\alpha)v^{\prime}(z)=0$. This implies
$zv^{\prime\prime}(z)+(1-z)v^{\prime}(z)=0,$
which is the confluent hypergeometric differential equation $zv^{\prime\prime}(z)+(c-z)v^{\prime}(z)-av(z)=0$ with coefficients $a$ and $c$ satisfying $a=0$ and $c=1$. This can be rewritten as $\frac{zv^{\prime\prime}(z)+v^{\prime}(z)}{zv^{\prime}(z)}=1$, so that $\frac{{\rm d}\ln(zv^{\prime}(z))}{{\rm d}z}=1$. The differential equation yields $\ln(zv^{\prime}(z))=z+c_{1}$, and thus $v^{\prime}(z)={\rm e}^{c_{1}}\frac{{\rm e}^{z}}{z}$, where $c_{1}$ is a constant of integration. Integrating the foregoing equation yields
$\int_{c_{2}}^{z}v^{\prime}(t){\rm d}t={\rm e}^{c_{1}}\int_{c_{2}}^{z}\frac{{\rm e}^{t}}{t}{\rm d}t,$
@@
$\int_{c_{2}}^{z}v^{\prime}(t){\rm d}t=\int_{-c_{2}}^{-z}w^{\prime}(s){\rm d}s=-{\rm e}^{c_{1}}\int_{-c_{2}}^{-z}\frac{{\rm e}^{-s}}{-s}{\rm d}s={\rm e}^{c_{1}}\int_{-z}^{-c_{2}}\frac{{\rm e}^{-s}}{-s}{\rm d}s={\rm e}^{c_{1}}\left(\int_{-z}^{\infty}\frac{{\rm e}^{-s}}{-s}{\rm d}s-\int_{-c_{2}}^{\infty}\frac{{\rm e}^{-s}}{-s}{\rm d}s\right).$
Let ${\rm Ei}(z)=-\int_{-z}^{\infty}\frac{{\rm e}^{-s}}{s}{\rm d}s$ denote the exponential integral. We then have $w(-z)-w(-c_{2})=v(z)-v(c_{2})={\rm e}^{c_{1}}[{\rm Ei}(z)-{\rm Ei}(c_{2})]$, which can be rewritten as $v(z)={\rm e}^{c_{1}}{\rm Ei}(z)+v(c_{2})-{\rm e}^{c_{1}}{\rm Ei}(c_{2})$. Setting $C=v(c_{2})-{\rm e}^{c_{1}}{\rm Ei}(c_{2})$ and $K={\rm e}^{c_{1}}$, and recalling that $v(z)=u(q)$ and $z=-\alpha q$, we obtain
@@
Appendix A.6 Proof of Corollary 2
@@
We thus obtain (15).
Equation (16) can be obtained as follows. Recall that $_{2}F_{1}\bigl(1,1;1+\frac{1}{\sigma};\frac{\gamma}{q}\bigr)$ in (15) is defined for $\lvert\frac{\gamma}{q}\rvert<1$. Thus, there are two cases depending on the sign of $\gamma$. [^38] Strictly speaking, there is another case with $\gamma=0$. Since $_{2}F_{1}\bigl(1,1;1+\frac{1}{\sigma};\frac{\gamma}{q}\bigr)=1$ for $\gamma=0$, we do not have to replace $_{2}F_{1}\bigl(1,1;1+\frac{1}{\sigma};\frac{\gamma}{q}\bigr)$ in (15) with the integral in (16). However, we can let $\gamma\to 0$ in (16) to obtain $u(q)=\kappa+\frac{\widetilde{\beta}\sigma}{\sigma-1}q^{\frac{\sigma-1}{\sigma}}$ for all $q\in[0,\infty)$, which is equivalent to (15) evaluated at $\gamma=0$. First, when $\gamma>0$, $\lvert\frac{\gamma}{q}\rvert<1$ implies $q>\gamma$, which is not restrictive because $u^{\prime}(q)=\widetilde{\beta}(q-\gamma)^{\frac{\sigma-1}{\sigma}}q^{-1}>0$ and $\widetilde{\beta}>0$ imply $q>\gamma$. Second, when $\gamma<0$, $\lvert\frac{\gamma}{q}\rvert<1$ implies $q>-\gamma$, which appears to be restrictive since (15) is not defined for $q\in(0,-\gamma]$. However, since $1+\frac{1}{\sigma}>1>0$ holds in Corollary 2, we can extend $_{2}F_{1}\bigl(1,1;1+\frac{1}{\sigma};\frac{\gamma}{q}\bigr)$ in (15) to $q\in(0,-\gamma]$ by analytic continuation (see, e.g., (9.1.6) in Lebedev, 1965, where $\lvert{\rm arg}(1-z)\rvert=\lvert{\rm arg}(1-\frac{\gamma}{q})\rvert=0$ since $1-\frac{\gamma}{q}>0$). Thus, in both cases, we can use the integral representation $_{2}F_{1}\bigl(1,1;1+\frac{1}{\sigma};\frac{\gamma}{q}\bigr)=\frac{\Gamma(1+\frac{1}{\sigma})}{\Gamma(1)\Gamma(1+\frac{1}{\sigma}-1)}\int_{0}^{1}(1-t)^{\frac{1}{\sigma}-1}\bigl(1-\frac{t\gamma}{q}\bigr)^{-1}{\rm d}t$, where $\Gamma(\cdot)$ is the gamma function. Using the recurrence relation $\Gamma(1+\frac{1}{\sigma})=\frac{1}{\sigma}\Gamma(\frac{1}{\sigma})$ and $\Gamma(1)=1$, the foregoing expression reduces to $_{2}F_{1}\bigl(1,1;1+\frac{1}{\sigma};\frac{\gamma}{q}\bigr)=\frac{1}{\sigma}\int_{0}^{1}(1-t)^{\frac{1}{\sigma}-1}\bigl(1-\frac{t\gamma}{q}\bigr)^{-1}{\rm d}t$. Hence, we can rewrite (15) as (16). $\square$
Proofs
Appendix B.1 Proof of Proposition 3
In the main text, we have already analyzed the case where either $\alpha=0$ or $\beta=\frac{1}{\sigma}$ holds. Thus, in this proof, we focus on the case where $\alpha\neq 0$ and $\beta\neq\frac{1}{\sigma}$.
@@ Appendix B.1 Proof of Proposition 3
We first derive the ranges of markups and quantities that satisfy the four inequality conditions, $q\geq 0$, $u^{\prime}>0$, $u^{\prime\prime}<0$, and $0<\mu\leq 1$. Table B.1 summarizes these ranges.
Table B.1: The ranges of markups and quantities satisfying $q\geq 0$, $u^{\prime}>0$, $u^{\prime\prime}<0$, and $0<\mu\leq 1$.
@@
$\Phi(q)\equiv\left\{\begin{array}[]{ll}\alpha^{2}(\sigma-1)\bigl[q+\frac{1-\beta}{\alpha(\sigma-1)}\bigr]^{2}-\frac{(1-\beta)(1-\beta\sigma)}{\sigma-1}&\mbox{for}\ \sigma\neq 1\\ (1-\beta)(2\alpha q+\beta)&\mbox{for}\ \sigma=1\end{array}\right.$. (B.1)
@@
$q^{\pm}=-\frac{1-\beta\pm\Delta}{\alpha(\sigma-1)},\quad\mbox{where}\quad\Delta\equiv\sqrt{(1-\beta)(1-\beta\sigma)}.$ (B.2)
@@
First, assume that $\sigma<1$. Since $1-\beta\sigma>0$ and $\beta\neq 1$ hold, $\Phi(q)=0$ has two real roots (B.2). Consider the subcase with $\beta\in(0,1)$. In that case, one of them is positive since $1-\beta+\Delta>0$, whereas the other is negative since $1-\beta-\Delta<0$. Furthermore, $\Phi$ is concave, so that it is positive for $0\leq q<-\frac{1-\beta+\Delta}{\alpha(\sigma-1)}$. We also know from Table B.1 that $q<-\frac{1-\beta}{\alpha(\sigma-1)}$ must hold. Since $\frac{1-\beta}{\alpha(\sigma-1)}<-\frac{1-\beta+\Delta}{\alpha(\sigma-1)}$ holds for $\sigma<1$, the SOC holds for all $0\leq q<-\frac{1-\beta}{\alpha(\sigma-1)}$ when $\beta\in(0,1)$. Consider the subcase with $\beta=0$. Then, the two real roots in (B.2) become $\frac{2}{\alpha(\sigma-1)}>0$ and 0. Since $\Phi$ is concave and $\frac{2}{\alpha(\sigma-1)}$ is greater than $\frac{1-\beta}{\alpha(\sigma-1)}$ in Table B.1 evaluated at $\beta=0$, the SOC holds for all $0<q<-\frac{1}{\alpha(\sigma-1)}$. As $q$ goes to zero, we know that $\frac{\Phi(q)}{(\alpha q+\beta)(\alpha\sigma q+1)}$ evaluated at $\beta=0$ goes to 2. Hence, the SOC holds for all $0\leq q<-\frac{1-\beta}{\alpha(\sigma-1)}$ when $\beta\in[0,1)$.
Second, assume that $\sigma=1$. Then, the SOC holds for all $q\geq 0$ since both $1-\beta$ and $2\alpha q+\beta$ in (B.1) are positive when $\beta\in(0,1)$ and since $\lim_{\sigma\to 1}\frac{\Phi(q)}{(\alpha q+\beta)(\alpha\sigma q+1)}=\frac{2}{\alpha q+1}>0$ when $\beta=0$. Hence, when $\sigma=1$ and $\beta\in[0,1)$, the SOC holds for all $0\leq q<\infty$ given in Table B.1 (recall that when $\sigma=1$, $0\leq q<-\frac{1-\beta}{\alpha(\sigma-1)}$ must be replaced with $0\leq q<\infty$).
Last, assume that $\sigma>1$. Since $1-\beta\sigma>0$ and $\beta\neq 1$ hold, $\Phi(q)=0$ has two real roots (B.2). Consider the subcase with $\beta\in(0,1)$. In that case, both are negative since $1-\beta+\Delta>0$ and $1-\beta-\Delta>0$. Furthermore, $\Phi$ is convex, so that it is positive for all $q\geq 0$. We also know from Table B.1 that there is no additional restriction on $q$, so that the SOC holds for all $0\leq q<\infty$ when $\beta\in(0,1)$. Consider the subcase with $\beta=0$. Then, the two real roots in (B.2) become $\frac{2}{\alpha(\sigma-1)}<0$ and 0. Since $\Phi$ is convex, the SOC holds for all $q>0$. As $q$ goes to zero, we know that $\frac{\Phi(q)}{(\alpha q+\beta)(\alpha\sigma q+1)}$ evaluated at $\beta=0$ goes to 2. Hence, the SOC holds for all $0\leq q<\infty$ when $\beta\in[0,1)$.
Since the range of $q$ in Table B.1 is unaffected by the SOC, that of $\frac{1}{\mu}$ in Table B.1 is also unaffected because $\mu=1-\frac{\alpha q+\beta}{\alpha\sigma q+1}$ is monotone with respect to $q$. Hence, the ranges for markups and quantities in Table B.1 remain valid in Table 3.
@@
If $\beta\in(0,1)$, then $\Phi(q)=0$ has no real roots since $1-\beta\sigma<0$. Since $\sigma>1$ and $\beta\not\in\{0,1\}$, $\Phi$ is convex and $\Phi(0)=\beta(1-\beta)>0$, which implies that $\Phi$ is positive for all $q\geq 0$. Hence, the SOC holds for all $q\geq 0$. Since the range of $q$ in Table B.1 is unaffected by the SOC, that of $\frac{1}{\mu}$ in Table B.1 is also unaffected because $\mu=1-\frac{\alpha q+\beta}{\alpha\sigma q+1}$ is monotone with respect to $q$.
If $\beta=1$, then $\Phi(q)=0$ yields $q=0$, which implies $\Phi(0)=0$. Since $\sigma>1$, $\Phi$ is convex, which implies that $\Phi$ is positive for all $q>0$. Hence, the SOC holds for all $q>0$. Since the range of $q$ in Table B.1 is unaffected by the SOC, that of $\frac{1}{\mu}$ in Table B.1 is also unaffected because $\mu=1-\frac{\alpha q+\beta}{\alpha\sigma q+1}$ is monotone with respect to $q$.
@@
First, assume that $\sigma<1$. Consider the subcase with $\beta\in(0,1)$. In that case, $\Phi(q)=0$ has two real roots (B.2). One of them is negative since $1-\beta+\Delta>0$, whereas the other root is positive since $1-\beta-\Delta<0$. Furthermore, $\Phi$ is concave, so that it is positive for $0\leq q<-\frac{1-\beta-\Delta}{\alpha(\sigma-1)}$. We also know from Table B.1 that $0\leq q<-\frac{\beta}{\alpha}$ must hold. Since $\frac{1-\beta-\Delta}{\alpha(\sigma-1)}<-\frac{\beta}{\alpha}$ holds, the second-order condition holds for all $0\leq q<-\frac{1-\beta-\Delta}{\alpha(\sigma-1)}$ but it does not hold for $\frac{1-\beta-\Delta}{\alpha(\sigma-1)}\leq q<-\frac{\beta}{\alpha}$. Since the bounds for $q$ are modified, we need to modify those for $\mu$. Noting that $\frac{\partial\mu}{\partial q}=\frac{\partial}{\partial q}\bigl(1-\frac{\alpha q+\beta}{\alpha\sigma q+1}\bigr)=-\frac{\alpha(1-\beta\sigma)}{(\alpha\sigma q+1)^{2}}>0$, we evaluate $\mu=1-\frac{\alpha q+\beta}{\alpha\sigma q+1}$ at $q=0$ and $q=-\frac{1-\beta-\Delta}{\alpha(\sigma-1)}<-\frac{\beta}{\alpha}$ to obtain $1-\beta\leq\mu<\frac{(1-\beta)\sigma+\Delta}{1+(1-\beta)\sigma}<1$. [^40] The last inequality can be obtained as follows. First, for $\frac{(1-\beta)\sigma+\Delta}{1+(1-\beta)\sigma}<1$ to hold, we require $\frac{-1+\Delta}{1+(1-\beta)\sigma}<0$. Second, the latter inequality requires that the numerator and the denominator must have opposite signs, which is indeed the case. Hence, we obtain the bounds for markups $1<\frac{1+(1-\beta)\sigma}{(1-\beta)\sigma+\Delta}<\frac{1}{\mu}\leq\frac{1}{1-\beta}$ for $\sigma<1$. Consider the subcase with $\beta=1$. Then, $\Phi$ attains the maximum $\Phi(q)=0$ at $q=0$. Since $\sigma<1$, $\Phi$ is concave, so that the second-order condition does not hold for $q\geq 0$, thus excluding the second line in Case 2a in Table B.1.